Can you help me solve (x=-2y+1)
(x=y-5)

Answer by sdmmadam@yahoo.com(530)

Can you help me solve (x=-2y+1)
(x=y-5)
x=-2y+1 ----(1)
x= y-5 ----(2)
(1)-(2) gives
(x-x)=[(-2y)-(y)]+[(1)-(-5)] (observing that the coefficient of x in both the equations is the same 1 and hence by subtraction the x variable is eliminated)
0= -3y+ 6
3y = 6
y = 6/3 = 2
y=2 in (1) gives x = -2y+1 = -2X(2)+1 =-4+1 = -3
Answer:x=-3 and y = 2
Verification: Putting x=-3 and y = 2 in the other equation x= y-5 ----(2)
LHS= x = -3 and RHS= y-5 = 2-5 =-3 = LHS.
Therefore our values are correct.